The ship left San Francisco at roughly 9:00 a.m., 20 September, 1996. Fresh northerly headwinds were encountered while sailing to Site 1023, which slowed the vessel's speed occasionally to less than 9 kt. Conditions continued to moderate throughout the transit, and speeds improved to 11.0 kt on the final day. A standard advanced hydraulic piston core/extended core barrel (APC/XCB) bottom-hole assembly (BHA) was deployed without a nonmagnetic drill collar or lockable float valve because core orientation and wireline logging were not scheduled for any of the HT sites. Rig floor operations were slowed on the initial pipe trip for several reasons. This was the initial pipe trip of the leg, and the stands of drill pipe had to be measured (strapped) and adequate internal diameter verified (rabbited). A drill string wiper plug (pig) was pumped to remove any loose scale residing in the pipe. In addition, a new driller and assistant driller were being trained. Operations went exceptionally well with no major problems or incidents. The mudline was verified at 5.2 m higher than indicated by the precision depth recorder (PDR). Hole conditions were excellent and no drilling mud was required. Seawater was the only circulation fluid used.
After releasing and recovering the beacon at Site 1023 (HT-2A) the ship was moved in dynamic positioning (DP) mode using differential Global Positioning System (dGPS) to the location coordinates for Site 1024. A single piston core was taken only to define mudline for the jet-in test. The mudline was verified to be 7.1 m higher than the PDR indicated. A jet-in test was conducted to 38.5 mbsf in 55.5 min, using up to 22 spm. This information was used to determine the amount of 16" casing (conductor pipe) to be deployed with the reentry cone on Hole 1024C.
The ship was offset in DP mode 500 m to the south-southwest before spudding Hole 1024B. The hole was located as far as possible from the coordinates of the primary CORK hole. The move was made with the drill pipe suspended on knobby drilling joints through the guide horn and the bit a minimum 100 m above the relatively flat seafloor. Based on the APC spud core, the mudline was verified 7.5 m above that indicated by the PDR. Hole conditions were again excellent and no drilling mud was required. The hole was displaced with weighted (10.5 lb/gal) gel mud after reaching basement depth. Operations went exceptionally well with no major problems or incidents..
After the beacon at Site 1024 was released and recovered, the ship was moved in DP mode using DGPS to the coordinates for Site 1025. A single piston core was taken to define the mudline for the jet-in test. The mudline was verified to be 6.2 m higher than the PDR indicated. A jet-in test was conducted to 39.7 mbsf using up to 60 spm and completed in 3 hrs and 18 min. This information was used to determine the amount of 16" casing (conductor pipe) to be deployed with the reentry cone on Hole 1025C.
The ship was offset in DP mode 500 m to the south-southwest before spudding Hole 1025B. The hole was located as far as possible from the coordinates of the primary CORK hole. The move was made with the drill pipe suspended on knobby drilling joints through the guide horn and the bit at a minimum 100 m above the relatively flat seafloor. Based on the APC spud core, mudline was verified 3.9 m above that indicated by the PDR. Hole conditions remained excellent and no drilling mud was required. The hole was displaced with weighted (10.5 lb/gal) gel mud after reaching basement depth. Operations went exceptionally well with no major problems or incidents.
Prior to spudding Hole 1025C the ship was moved in DP mode back 500 m to the original coordinates for Hole 1025A. A reentry cone and 16" conductor pipe were deployed and washed into the seafloor without incident. The 16" casing shoe was placed at 2657.4 meters below rig floor (mbrf) (40.2 mbsf). The Dril-Quip 16" running tool was released from the reentry cone/casing. The running tool was dressed with the CADA option allowing the 14-3/4" tricone drill bit to be advanced downhole. A 14-3/4" hole for the 10-3/4" surface casing string was drilled to 106.1 mbsf (5.1 m into basement). Two meters of hard fill were found in the hole after picking up the drill string for a connection. Drilling progress slowed noticibly at that point, and it was decided not to pursue deepening the hole farther. The hole was flushed with a 30 bbl sweep of sepiolite mud, and the bit was pulled to 2644.0 mbrf (up inside the 16" conductor casing). To allow time for any fill to settle in the hole, the top drive was set back and the drilling line was slipped and cut. Approximately 2.5 hr later the second half of the wiper trip was begun. The pipe was run to the original total depth (TD) of 2723.3 mbrf (106.1 mbsf) without any resistance. Encouraged by the apparent stability, we left the hole filled with seawater and the drill string was tripped back to the surface. A 98.8-m-long string of 40.5 lb/ft 10-3/4" K-55 casing was then made up and run to bottom. Reentry took 1 hr partially because of obscured vision from sediment clouds emenating from the reentry cone. Once the hole was reentered, the casing went smoothly downhole without any problem. Approximately 4 m short of landing the casing hanger, the drill pipe was hung off at the rotary table. The top drive was picked up along with the TIW subsea-release cementing manifold and a 30 ft knobby drilling joint. With the heave compensator activated, the 10-3/4" casing hanger was landed and latched. A 10,000 lb tension was taken on the hanger to verify proper latch-in. With the 10-3/4" casing shoe placed at 2718.7 mbrf (101.5 mbsf), 44 bbl of 15.8-ppg class G neat cement were mixed and displaced around the casing shoe. With a 5,000 lb tension applied, the drillstring was rotated 3-1/2 turns to the right. After a few minutes of working the pipe and releasing the torque buildup, the Dril-Quip running tool spun freely. A 12,000 lb weight loss indicated that the casing was free of the drill string and the pipe was tripped back to the surface. The drill floor was secured for transit, the positioning beacon was commanded off, and the vessel then got underway for proposed Site PP 5A (Site 1026). Operations moved to the RB area to allow the cement ample time to cure and, thus, to increase the chance for a better seal. This strategy was used for all the cased holes in which a CORK was installed, which enabled us to use the time spent waiting for cement to cure to complete coring and casing at nearby holes.
The 3.5 hr transit to Site 1026 was brief and uneventful. Once the beacon was deployed, the pipe was run to bottom and APC coring commenced. The first core recovered 5.41 m, which established the mudline 3.3 m higher than the PDR indicated. The hole was cored to APC refusal (as defined in this hole by failure to fully stroke). Adara temperature measurements were taken on every other core beginning with Core 168-1026A-4H. Drilling operations were routine with no major problems or incidents. Seawater was the only circulation fluid used during the coring operation. The hole was displaced with heavy mud (10.5 lb/gal) prior to abandonment. After clearing the seafloor with the drill bit, a jet-in test was conducted to determine the amount of 16" casing (conductor pipe) to be used with the reentry cone for this site. The beacon was left on bottom and turned on, because the site was to be reoccupied relatively soon.
The drill ship was offset 1.2 nm to Site 1027 in DP mode with the drill string suspended in the same manner as on the other DP moves this leg. Based on earlier mudline depths, the bit was positioned 6.4 m above the PDR depth for the first spud attempt. Sediment was penetrated, but the first core barrel recovered was empty; the core had fallen out of the liner, and, thus, it was impossible to identify a mudline depth. A second attempt at spudding was made from the same depth, and this time a full core (9.92 m) was recovered. The hole was abandoned again for lack of a positive mudline identification. The hole was logged as an official hole (1027A) because the core material was curated and kept for further analysis by the shipboard geochemists.
After a successful mudline core was recovered, coring in Hole 1027B continued through Core 168-1027B-62X. Basalt pebbles were first identified in Core 168-1027B-60X and became more predominant with each successive core. The hole was terminated at a depth of 3246.2 mbrf (577.9 mbsf) when it was decided not to spend any more time searching for "absolute" basement. Sepiolite mud sweeps of 25 bbl each were circulated approximately every 40-60 m to aid in hole cleaning. Despite high sand content in the upper formation, the hole remained stable to TD. There was no indication of fill after pipe connections and no increase in pump pressure or drilling torque as the hole proceeded. Once coring operations were concluded 25 bbl of 15.8-ppg cement were placed in the hole and the remainder was displaced with heavy mud.
A total of four Adara temperature measurements were taken in Hole 1027B. The DVTP was deployed 18 times. Tool 1 failed during recovery (after Core 168-1027B-40X). The data were ultimately recovered but are not a good measurement.
The ship was offset in DDP mode 50 m south-southwest and a jet-in test was conducted to a depth of 40.0 mbsf. A total of 1 hr 52 min was required for the test at a maximum of 50 spm and 10,000 lb weight. After recovering the drill string, a reentry cone with 16" casing was made up and deployed. The cone landed at the seafloor after jetting the casing shoe for 2 hr 39 min at up to 52 spm and using 20,000 lb weight. The 16" casing shoe was placed at 2704.9 mbrf (37.6 mbsf). The 16" Dril-Quip running tool was released and the drill string was tripped back to the ship. A 14-3/4" tricone bit and drilling BHA was then made up and tripped to bottom. The cone was reentered at 0150 hr on 8 July 96. A 14-3/4" hole was control-drilled to a depth of 3252.1 mbrf (584.8 mbsf) or 9.3 m into basement (575.5 mbsf). A wiper trip was made to the 16" casing shoe and back with minor overpull and drill pipe drag in spots. An apparent ledge was identified approximately 1 m into basement, and the bit was worked through this area several times. Tight hole and light fill characterized the lower 15 m. Without setting back the top drive, a short (six stand) wiper trip was made, and the hole was circulated clean and then displaced with sepiolite mud. The drill string was recovered and 41 joints of 10-3/4" casing were made-up and deployed using the Dril-Quip running tool. After we returned to the seafloor with the casing and searched for several hours, it became apparent that the reentry cone had sunk significantly below the mudline. The cone/16" casing became undermined during the drilling of the 14-3/4" hole. It is theorized that a combination of circulation and vibration fluidized the heavily sand-laden sediments in the upper portion of the hole, allowing the cone to sink 8.3 m below the mudline. Without seeing the cone, the hole was reentered using sonar and the 10-3/4" casing string was deployed until it landed firmly at a depth of 3252.05 mbrf, essentially the hole TD. The string was picked up, to no avail, three times in an attempt to verify that the casing hanger had engaged. The casing shoe was set down one last time while spotting 56 bbl of neat cement around the casing shoe. When the drill string was raised the final time, to everyone's surprise the reentry cone and 16" casing string came up with it. The reentry cone mudskirt was placed 2 m below the original mudline at a depth of 2669.3 mbrf. This positioned the 16" and 10-3/4" casing shoes at 2704.9 mbrf (38.6 mbsf) and 3245.7 mbrf (578.4 mbsf), respectively. The 10-3/4" shoe was cemented 2.9 m into basement and the entire rathole below the shoe was also filled with cement. Drill string tension was maintained for 12 hr to support the weight of the cone and casing strings while the cement hardened to an acceptable state. The 10-3/4" Dril-Quip running tool disengaged without incident. The cone was left stable and the surface sediments appeared to have backfilled into the cavity up to the bottom of the reentry cone mudskirt. Once the drill string was recovered to a safe depth, the positioning beacon was commanded off and the vessel began moving in DP mode back to Site 1026 (PP-5A).
The ship arrived back at Site 1026 before the drill string was completely recovered. The positioning beacon was commanded back on. The Dril-Quip 10-3/4" running tool was detorqued and the drilling line was slipped and cut before preparations began for deployment of the third reentry cone of the leg. After three joints of 16" casing were made up, the reentry cone and casing were deployed to the seafloor. Weather conditions were beginning to deteriorate at this point, so one stand of drill collars above the running tool was left out to facilitate getting the reentry cone through the surge zone of the moon pool as quickly as possible. The cone/casing were deployed to the seafloor. The jetting process required a total of 4 hr 9 min, with circulation rates of up to 51 spm and weights of 25,000 lb. The cone base plate landed solidly at the appropriate mudline depth (2669.1 mbrf) and the 16" Dril-Quip running tool was released routinely. After the drill string was recovered, the 16" Dril-Quip running tool was detorqued at the rig floor and the dart was loaded in the cementing swivel in preparation for later cementing operations. A 14-3/4" BHA was made in just a few minutes and the bit was run to the bottom of the jetted 16" casing. A total of 7.5 hr was required to drill the 14-3/4" hole to a firm basement depth of 2916.2 mbrf (247.1 mbsf). Another 3.25 hr of basement drilling brought the hole to a total depth of 2925.1 mbrf (256.0 mbsf). The hole was flushed with 40 bbl of sepiolite, and a wiper trip was made up into the 16" casing shoe. To allow time for the hole to stabilize and/or for cuttings to settle out, the subsea camera was deployed to inspect the condition of the reentry cone. The inspection indicated that the cone was in good condition, resting at the seafloor as designed. After recovering the camera, the pipe was run back to TD, where 20 m of loose fill was identified on bottom. The fill was circulated out to within 1 m of total depth without requiring rotation of the drill string. Another 40 bbl sepiolite mud sweep was pumped followed by a seawater spacer and enough sepiolite to displace the hole. The drill string was brought back to the surface and preparations were begun for making up 18 joints of 10-3/4" casing. This process was slowed somewhat by significant vessel motion, but was still completed safely and in timely fashion. The 10-3/4" casing string was deployed to the seafloor and, in spite of 1-2 ft heave conditions, the reentry was made in 45 min. The casing string was run to bottom without any resistance and latch-in was verified with 10,000-lbs overpull. The sepiolite in the hole was displaced out and 49 bbl of 15.8-ppg class G neat cement were pumped downhole. The subsea release (SSR) plug was released at 2800 psi and the cementing dart was landed at 1000 psi. Attempts to release the Dril-Quip running tool were unsuccessful, however, when the casing hanger and casing string appeared to be rotating with the running tool. At 1700 hr it was decided to wait and let the cement harden rather than risk compromising the cement bond and seal with the formation. After waiting on cement for 4 hr, the shipboard cement samples were checked and it was decided that another 2-4 hr should be allowed before attempting to put torque into the running tool. At 0300 hr on 15 July 96, another attempt was made at releasing the running tool. This time the running tool released properly and the drill string was recovered aboard ship. Once secured, the vessel got underway in DP mode back to reentry Site1027.
While the ship was moving to Hole 1027C, an attempt to detorque the 10-3/4" Dril-Quip running tool was made at the rig floor. This was only partially successful because of the large amount of fine sand that had infiltrated the mechanism.
The drill string with an RCB BHA was run in the hole with a center bit installed for drilling out the rubber cementing dart, casing shoe, and wiper plug. Reentry was made swiftly (<15 min) and the bit was run to bottom. The cement was tagged at 3234.7 mbrf (567.4 mbsf) and it took a total of 5 hr to completely drill out to the original TD of 3252.1 mbrf (584.8 mbsf). Continuous RCB coring continued from that depth through interbedded basalt flow units, mudstones, and highly altered basal sediments. Some overpull and torque were experienced during coring, but, in general, hole stability was good. No drilling mud was used while coring below the 10-3/4" casing shoe. The hole was eventually terminated in altered pillow basalt. The subsea TV was deployed on the VIT sleeve to inspect the reentry cone for sediment and cuttings. The cone was found clean and, thus, no jetting was required on the way out of the hole. A short wiper trip back to bottom indicated that the hole had 2.0 m of fill. This was circulated out and the drill bit was pulled clear of the reentry cone. Once secured, the ship was moved in DP mode back to Hole 1026B.
While the VIT sleeve was being deployed at Hole 1026B the subsea TV system went down, requiring 45 min to repair. The problem was traced to power supply filters installed in surface equipment in the DP control room. With the TV problem corrected, the reentry took place in short order (<15 min). The bit was run to bottom, tagging the cement at 2910.0 mbrf (240.9 mbsf). It took 5.5 hr to drill out the cement and shoe hardware, including recovery of the center bit. Continuous RCB coring proceeded to a depth of 2964.3 mbrf (295.2 mbsf). Coring was plagued with major hole problems. High torque, overpull, and high pump pressure were common. Recovery was extremely poor in the fresh, highly fractured basalt. Recovery consisted primarily of rubble with few cored pieces. Several mud sweeps totaling 160 bbl were circulated while attempting to clean and stabilize the hole. The more the hole was fought, the more unstable it became. Concerns over permanently sticking the drill string eventually led to the termination of coring. The last tag with drill pipe, using no rotation or circulation, indicated a final hole TD of 2939.0 mbrf (269.9 mbsf), or 25.3 m shallower than the maximum depth achieved during coring. The drill string was recovered and packer operations commenced. All pressure piping and manifolds were tested for leakage, and a wiper plug was pumped downhole to ensure that any loose rust would not infiltrate the packer setting go-devil.
After a pipe trip and reentry, the TAM (manufacturer) packer was run to the bottom of the 10 3/4" casing string. Prior to setting the packer, the wireline was run to bottom with the DVTP to get a bottom-hole temperature and to check the hole TD. The depth check indicated that the hole had continued to deteriorate; the fill level was now up to 2929.0 mbrf, or only 4 m below the depth of the rathole. A temperature measurement at TD indicated that the hole seemed to be losing fluid at a slow rate. Once inside the casing, the TAM packer was set at 2878.1 mbrf. Two slug tests and two injection tests were conducted. Early results indicated that the formation was slightly underpressured. Possible fluid loss through the formation or past the casing cement was also suspected. Upon completion of the packer experiments, the drill string was pulled clear of the reentry cone. The pipe was secured and the ship moved in DP mode back to Hole 1027C for additional packer work.
After moving back to Hole 1027C (PP-4A) in DP mode, the hole was reentered and the pipe was run to one stand above the 10-3/4" casing shoe. At that point the wireline sinker bars were deployed with the DVTP tool. A depth check of the hole indicated that the hole depth had remained stable and no further fill had entered. A bottom-hole temperature measurement was also taken at that time. The TAM packer was set twice. The first set was inside the casing at 3228.0 mbrf. Three slug tests and two injection tests were conducted on the first set. The second packer set was in open hole at 3258.0 mbrf (590.7 mbsf). Two slug tests and three injection tests were conducted on the second packer set. Early results indicated that the formation may again be slightly underpressured. The testing went exceptionally well and the data appear to be of good quality. After completion of packer experiments, the drill string was recovered and preparations were begun in calm seas for the first CORK deployment of the leg. Deployment of the CORK, data logger, and osmotic sampler systems went well and there were few problems. The CORK running tool was difficult to align with the CORK body because of their massive sizes. Once this was done and the mating surfaces were coated with pipe dope, the tool slid on easily. The thermistor string was lengthened by 15 m to take advantage of the extra hole depth. While the drill string was being recovered, the positioning beacon was released/recovered and the ship began moving in DP mode back to Hole 1026B (PP-5A) for final CORK operations at that site.
Before the drill string was fully recovered, the move in DP mode to Hole 1026B was completed. The CORK running tool was set aside and preparations were begun to deploy a drill-in liner assembly made up of a modified mechanical bit release (MBR) and seven joints of junk 5" drill pipe. The pin connection was cut off the MBR bit disconnect. The box end of the top joint of 5" drill pipe was then slipped on and welded. Stabilizer pads were welded onto the MBR body and turned to the drift diameter of the 10-3/4" casing. This assembly was run in the hole and the cone was reentered. The level of fill in the hole was tagged with the pin end of the drill-in liner at a depth of 2929.0 mbrf (259.9 mbsf). This was 35.3 m above the original hole TD and only 12.8 m into basement (2916.2 mbrf). It was hoped that by drilling/washing in the drill string liner a conduit would be provided farther into basement for the thermistor string. Because of the poor hole conditions, it was considered unlikely that the hole would ever be deepened in the future. After multiple attempts, the liner was eventually worked down to a depth of 2955.0 mbrf (285.9 mbsf) using very slow rpm (5) and very low circulation (20 spm). Drilling torque was 75-100 amps. Heavy mud was spotted in the drill string to keep the flow moving in a downward direction during connections. We tried to get the liner in place with the least disturbance or vertical displacement of the fill up the annulus outside the liner. It was feared that if they were highly fluidized, cuttings could flow back into the liner from the bottom or if displaced high enough would spill over the top of the liner. The rotary shifting tool and sinker bars were installed during the last connection, making the release operation relatively straight forward. Once the MBR was released, the annulus was circulated clean and the heavy mud was displaced out of the annulus using seawater. The shifting tool was recovered and the sinker bars were run back to bottom to check the depth of open hole. On the first tag, 4.0 m of fill was found inside the drill pipe liner at a depth of 2951.0 mbrf. After waiting 1 hr for cuttings to settle out further, the fill was again checked. This time the fill had reached a depth of 2949.0 mbrf, or 7 m up inside the liner. The final hole depth available for thermistor deployment was left at 31.8 m into basement. Nearly 20 m of open hole had been regained by this effort. The top of the liner was placed at 2887.0 mbrf, or 30.6 m above the 10-3/4" casing shoe. By the time the drill string was recovered, weather conditions were marginal for a CORK deployment, and the forecast indicated that weather and sea-state conditions would likely continue to deteriorate over the next 12-18 hr. As a result, the CORK deployment was deferred until later in the leg. The ship was secured and departed for Hole 1024C, where the last cone/casing deployment of the leg was to be made.
Because of poor weather conditions, we had to delay installation of a CORK at Site 1026. Therefore, we decided to move back to Site 1024 and set what had become the fourth reentry cone of the leg, in preparation for installing a CORK. After preparation, a cone with 16" conductor pipe was run to bottom. The jetting operation required a maximum of 46 spm and was completed in 2 hr 44 min. With the reentry cone landed at the mudline, the 16" casing shoe was placed at 2662.5 mbrf (39.2 mbsf). The Dril-Quip 16" running tool was easily released from the reentry cone and casing. The drill string was recovered, the 16" Dril-Quip running tool was detorqued at the rig floor, and a 14-3/4" drilling assembly with a tricone bit was made up. The assembly was lowered to the seafloor, and after deploying the VIT, a routine reentry was made in 30 min. The drilling rate to basement was fairly rapid at first. Drilling continued at an average rate of penetration (ROP) of 30 m/hr until a depth of 2775.0 mbrf (151.7 mbsf) was reached. After the drilling break, the ROP slowed to 1-2 m/hr until TD was reached at a depth of 2797.0 mbrf (173.7 mbsf), 22.0 m into hard rock. In the lower part of the hole, the drilling was characterized by high torque and some pipe sticking, but only on bottom. The drilling parameters rapidly returned to normal once the bit was lifted off bottom. At the conclusion of drilling, the hole was circulated with seawater and a wiper trip was made to the 16" casing shoe. During the return trip, fill was tagged 16 m off bottom. This was easily circulated away until firm fill was reached 7.6 m above the original TD. An additional 2.0 m was washed away and then a 40-bbl sepiolite mud sweep was pumped. After a short trip, fill was still evident 10 m off bottom. The top drive was used to circulate the fill out until hard fill was reached 3.0 m above TD. At this point another 50-bbl sepiolite pill was circulated followed by a 50-bbl seawater spacer and 140 bbl of sepiolite mud to displace the hole. The drilling assembly was recovered and preparations were begun for making up and running the 10-3/4" surface casing string. It took only 3.5 hr to make up 12 joints of K-55 40.5-lb/ft surface casing including making up the casing hanger, Dril Quip running tool, etc. After the casing was run to the seafloor, the top drive and TIW cementing manifold were picked up and within 45 min another reentry was made. The casing hanger was landed and latch-in confirmed with 15,000-lb overpull. A total of 61.7 bbl of 15.8-ppg class G neat cement was mixed and displaced downhole; however, there was no indication that the dart landed or that the wiper plug released properly. Although the Dril-Quip running tool appeared to rotate downhole the requisite number of turns (3.5 to the right) the tool would not come free from the hanger. After waiting nearly 6 hr for the cement to partially harden, the pipe was merely raised and the tool came free. Apparently the tool did rotate properly, because the buoyancy effect of the casing in the cement put the release ring in compression, preventing the ring from moving inward to release. A higher overpull applied earlier probably would have released the tool at that time. The drill string was tripped back to the drill ship, ending the present round of operations at Hole 1024C. This hole was to be reoccupied later for RCB coring in basement, packer experiments, and CORK setting.
The drill ship was offset in DP mode back to Hole 1025C while recovering the casing running BHA from Hole 1024C (HT-3A). The 10-3/4" Dril-Quip running tool was detorqued at the rig floor and an RCB BHA with mechanical drilling jars was made up and run in the hole. The reentry cone was reentered 15 min after reaching the seafloor, and the drilling assembly was run to bottom. The top of the cement was tagged at a depth of 2706.0 mbrf (88.8 mbsf). Just under 6 hr was required to drill out the cement in the casing, casing shoe, and cement beneath the shoe. The center bit was recovered on the wireline and RCB coring operations commenced. Continuous RCB coring in igneous basement progressed through Core 168-1025C-5R to a total depth of 2764.4 mbrf (147.2 mbsf) with moderate recovery. The rate of penetration varied significantly as various units within basement were penetrated. Drilling was slow in massive layers; some zones were encountered where the bottom seemed to fall out and rapid penetration was achieved for several meters. Hole conditions were generally good, although some pipe sticking and hole trouble became apparent after advancing Core 168-1025C-4R to a depth of 2754.8 mbrf (137.6 mbsf). The tight spots were worked repeatedly with the drill pipe, and 30 bbl sepiolite mud sweeps were pumped after Cores 168-1025C-4R and -5R. A wiper trip to 108.9 mbsf was made, and on the return trip ledges in the hole were identified at 2725.0 mbrf (107.8 mbsf) and 2746.0 mbrf (128.8 mbsf). Fill was tagged at 2749.0 mbrf, 15.4 m off bottom. Another 30-bbl sepiolite mud sweep was pumped while washing back to TD. On the next short trip, fill was tagged at 2756.0 mbrf, or 8.4 m off bottom. The fill was circulated out one last time, and a 30-bbl sepiolite mud sweep was pumped. The drill string was recovered and preparations for deployment of the TAM packer were initiated. Once reentry was accomplished with the TAM packer (30 min), the BHA was run to the packer setting point inside the second casing joint from the bottom. A wireline depth check of TD was made before beginning packer operations to determine the amount of hole available for thermistor emplacement. The DVTP was deployed and measurements taken at 5.0-m intervals from the 10-3/4" casing shoe to TD during the first depth-check wireline run. A temperature profile of the hole was obtained. Hole depth was measured at 2754.0 mbrf (136.8 mbsf), or 35.8 m into basement. Several unsuccessful attempts were made to set the TAM packer. Multiple wireline runs were made with the first setting go-devil, and another go-devil was used as well. Nothing was wrong with either go-devil yet the packer would not inflate. Another depth check of the hole was made using the wireline while meeting with rig personnel, co-chiefs, and packer scientists. It was decided that further attempts to perform packer operations would be abandoned in lieu of proceeding with setting the CORK assembly. Two compelling factors affected the decision. The current weather window was ideal for CORK setting and the second wireline depth check indicated another possible loss of 2.0 m of hole (TD 2752.0 mbrf). The measurements were made about 6 hr apart and could have been partially influenced by the 2.0-m tide that was apparent for most of the leg; however, there was continued concern about additional loss of available hole for thermistor deployment. The drill string was tripped back to the rig floor and preparations for setting the second CORK were begun.
Upon inspection, no mechanical problems were found with the packer assembly. A new control sleeve was installed after the last deployment. The new sleeve apparently fit somewhat tighter than the old worn one. With little or no weight below the packer, the new control sleeve apparently stopped approximately 1" from being fully extended in the open position. If the control sleeve is not fully open, inflation cannot take place and setting the packer becomes impossible.
Obsolete 9-1/2" drill collars (three each) were used for CORK stinger material. These were too large for the iron roughneck to handle, so conventional rig tongs were used for make up. Rig-up and deployment of the CORK assembly went well without major problems. The CORK was tripped to the seafloor and reentry was made within 10 min. The stinger was positioned inside the reentry cone without landing the CORK and the thermistor string/osmotic pump assembly was deployed down the drill string. As with the first CORK deployment at Site 1027, the dual grip lifting technique was used. This system is less than ideal, but it worked well enough for the relatively short thermistor string. Within 1 hr the thermistor was deployed inside the drill pipe and ready for running to bottom. A final electronic check was made on the data logger located at the top of the thermistor string and the assembly was run to bottom on the wireline using the setting tool of the datalogger. Several landing attempts were made to seat the seals before the release tool would pressure up. The shear pin in the overshot apparently sheared during the pressurization phase, because no overpull was experienced when recovering the wireline. The CORK setting go-devil was deployed after pulling the mousehole and picking up two additional joints of drill pipe. The CORK was landed and the setting go-devil pressured up. Latch engagement was verified with drill string overpull and the CORK platform was deployed. The subsea TV was deployed to verify that the installation was complete. The CORK running tool was released and the camera was recovered. The drill string was tripped back to the rig floor and the positioning beacon was recovered. During the trip out, the ship was offset in DP mode back to Hole 1024C. The hole was ended once all tools were recovered back aboard ship.
An RCB BHA was made up and run to the seafloor. The VIT was deployed during the pipe trip and a routine reentry was made within 15 min. The bit contacted the top of the cement column at a depth of 2777.2 mbrf (153.9 mbsf). The cement and float shoe were drilled out in short order, and the center bit was recovered. On this hole the cementing dart and wiper plug did not release. The dart was recovered, but it was not known whether the casing wiper plug fell into the reentry cone or outside the hole. The bit broke through to the rathole after only 1.5 hr. Once RCB coring was initiated, the bit advanced only 2.0 m before tight hole and torquing led to early retrieval of the core barrel. Recovery was minimal (0.12 m), but included in the core liner was a section of the float shoe including the entire float valve assembly. Apparently the inner portion of the float shoe had broken free during the center bit drilling phase and fallen to the bottom of the rat hole. This allowed the core barrel to swallow the float valve perfectly during the first coring attempt. The next 4 hr were spent fighting stuck pipe and hole trouble until the decision was made to abandon attempts at further penetration. Earlier in the leg, while drilling the 14-3/4" hole for the 10-3/4" surface casing, the drilling was also stopped because of bad hole conditions and stuck pipe. The hole was abandoned at that time at 22.0 m into basement and was left with 3.0 m of hard fill on bottom. Coring operations on this bit run were halted after one core and a total TD of 2797.0 mbrf (173.7 mbsf). The depth of open hole remained the same, at 2794.0 mbrf (169.7 mbsf). The bit was pulled up inside the 10-3/4" casing shoe and the second core barrel was recovered. This was not considered a core, because no advance was made and no core recovery was obtained. After tripping the drill string and prior to making up the TAM packer BHA, the mechanical drilling jars were tested at the rig floor. The jars were not only locked up from the debris but also had failed to hold circulation pressure. This undoubtedly contributed to the hole cleaning problems during the attempted coring, because full circulation was obviously not getting to the bit. Because identical drilling problems occurred while drilling the 14-3/4" hole without the drilling jars in the BHA, we did not believe that was the sole reason for the failure to advance the hole. The jars were cleaned up and prepared for shipment back for refurbishment. A reentry cleanout bit and a 10-ft drill collar pup joint were made up directly below the packer to aid in keeping the control sleeve fully extended. The packer BHA was run to bottom, but prior to reentry a wiper plug was pumped down the drill string. After another routine 15-min reentry, the circulation system was pressure tested and the pipe was lowered to a packer setting depth of 2775.3 mbrf (152.0 mbsf). The DVTP was run to bottom on the wireline for a preliminary depth check of the hole and to obtain a bottom-hole temperature. There was no problem setting the packer in this hole, and the packer testing went exceptionally well. After two successful injection and flow tests were completed, the packer was released and a final depth check was made with the wireline. Useable hole depth was confirmed as unchanged at 2794.0 mbrf, and this depth was used to calculate the final thermistor length for the hole. The pipe was retrieved back to the ship and preparations began for deploying the third CORK. The CORK deployment again went exceptionally well. The last of the obsolete 9-1/2" drill collars (three each) were made up as the CORK stinger and hung off. The CORK body was positioned in the skate and the running tool made up. The CORK assembly was made up to the CORK stinger and lowered so the final connection could be made with the CORK setting hose. As on earlier deployments, the hole number was painted on the CORK body identifying the installation for later visits by submersible or remote-operated vehicle (ROV). Once the safety blocks and packing protectors were removed, the CORK was ready for deployment. The tool was run to bottom, and within 15 min a reentry was made. Deployment of the osmotic sampler, thermistor string, and data logger proceeded without incident. A final electronic check on the data logger was conducted just prior to installing the running tool. The thermistor string was run to bottom and the data logger landed/latched into the CORK. A wireline overpull, before shearing the weakened overshot pin in the running tool, confirmed data logger latch-in. After retrieving the wireline, the CORK setting tool was deployed and pressured up. After bleeding off the pressure, the drill pipe was picked up, and a 10,000-lb overpull confirmed that the CORK was now latched in place. The VIT frame was deployed and the subsea TV was used to witness the release from the CORK installation. The setting go-devil was retrieved and the drill string was tripped back to the ship while the positioning beacon was released and recovered. Once the ship was secured, the hole was officially ended and the transit to Hole 1026C was begun.
After the completion of the CORK operations at Sites 1024 and 1025, the ship was secured and headed to Site 1026 to complete the fourth and last CORK installation and to core the lower part of the sediment section. During the 5-hr transit back, the ship was slowed to 6 kt to test a new seismic streamer. Once the test was completed, the ship resumed full speed. Hole 1026C was located approximately 1000 m from Hole 1026B, at N17ĦE, but was geologically considered part of the same sequence of holes at Site 1026. A beacon was dropped and an RCB BHA was made up and tripped to bottom. The RCB system was used because there was higher interest in recovering indurated sediment and basement samples rather than the softer overlying sediments. Because the upper 101.4 m of sediment had already been piston cored at Hole 1026A, the top 84.6 m was drilled with a center bit. Continuous RCB coring began at that point and continued to a depth of 248.2 mbsf. Recovery was poor until nearly 210 mbsf, where the sediments were finally indurated enough for the RCB system to be effective. The last three sediment cores achieved 90% recovery. Two subsequent basement cores in basalt rubble again failed to recover any "cored" material providing only pieces of rubble "rollers" that tended to jam the core catchers. Some torquing was evident in basement, but was not severe. One meter of fill was detected after making the connection before the last core. Once coring operations were terminated, the pipe was worked back to bottom and a 30-bbl sepiolite mud pill was circulated to try and clean the hole as much as possible. A wireline run was made to release the bit and the hole was plugged with 29.7 bbl of 15.8-ppg class G cement spotted in basement. The cement was required to reisolate basement hydrologically to avoid possible perturbations at the nearest CORK site at Hole 1026B. The bit was then pulled clear of the mudline, and the beacon was recovered to end Hole 1026C.
The ship was offset 1000 m back to Hole 1026B, and the hole reentered within 15 min after the rig floor was ready. The pipe was lowered to a depth of 2874.6 mbrf, and after the top drive and a 20-ft knobby joint were picked up, the DVTP was run in on the wireline. A temperature profile of the borehole fluids was recovered as a bonus while running to bottom for a final thermistor string depth check. Fill was tagged at a depth of 2947.0 mbrf. This was considered essentially the same as before, because it was within 1 m of the last measured depth (i.e., within the variation of a 2-m tide). The top drive was set back, and the driller began to trip the drill string back to the ship. Approximately half way into the pipe trip, the evaluation of the temperature data from the last DVTP deployment was completed. The record indicated an influx of warm formation fluids into the hole. This was seen as an excellent opportunity to attempt recovery of pristine basement fluids. Unfortunately, the WSTP tools would not fit through the bore of the CORK assembly. Therefore, the pipe trip was reversed and the driller began to run back to bottom. After deploying the subsea TV camera, another quick reentry was made and the pipe was run to a depth of 2720.8 mbrf. The top drive was again picked up and the WSTP was run into the hole with an Adara temperature measurement shoe made up to the bottom for redundancy. It was determined that the most likely path for fluid flow was up the annulus between the makeshift drill pipe liner installed earlier in the leg and the 10-3/4" casing. Therefore, the water sampler probe was positioned 8 m above the top of the liner. After allowing time for the sampler valve to open and close, a more detailed temperature profile was taken from the liner to the mudline. Measurements were taken every 10 m for 3 min until 20.9 mbsf. Then the measurements were taken every 2.0 m until 9.1 m above the mudline. When the tool was recovered, it was kept on the rig floor until the quality of the data could be confirmed. The tools were deployed a second time because the volume of fluid recovered was inadequate and the memories on both the Adara and WSTP temperature tools ran out before the mudline was reached. The second deployment recovered a large volume of high-concentration formation fluid and the temperature data were completed. The exercise was exceptionally successful, and all concerned agreed that the time was well spent and the data and samples recovered were extremely valuable. By the time the drill string was recovered, weather conditions had begun to deteriorate and the prognosis was for continued 25-30-kt winds over the next 12-24 hr. Because the sea state had already begun to build, it was deemed inadvisable to begin CORK operations. Once again, deployment of the CORK was deferred at Hole 1026B. The ship was secured and then departed for Site 1028.
The 31.0-nmi transit to Site 1028 was made at a speed of 8.4 kt. A couple of hours were spent at reduced speed (6.0 kt) to perform another test of the seismic streamer. Once on site, the beacon was released and an APC/XCB BHA was made up and run to the seafloor. APC coring proceeded to a depth of 108.2 mbsf where four Adara temperature measurements were taken. A maximum overpull of 50,000 lb was recorded with the Adara shoes and of 30,000 lb without any Adara shoes. XCB coring continued to a final TD of 131.5 mbsf. Basement was contacted at a depth of 130.5 mbsf in Core 168-1028A-15X and 1 hr was spent coring basalt with a hard formation cutting shoe. The DVTP was deployed twice in the XCB-cored interval. The weather moderated significantly, and within 1/2 hr of completing the pipe trip, the ship was underway for Hole 1026B to deploy the fourth and final CORK of the leg.
Return to Hole 1026B
After completion of Hole 1028A, the ship sailed a few hours in calm seas back to Hole 1026B. As on an earlier transit, speed was reduced to 6 kt for a brief period (2 hr) for a second test of the developmental seismic streamer. Once on site, the positioning beacon was turned on and preparations began quickly for deploying the CORK assembly. The entire CORK deployment operation, less the return pipe trip, was completed in only 14 hr. This includes the trip to the seafloor (2669.1 mbrf) with the CORK assembly, handling of the osmotic sampler, deployment of 277.5 m of thermistor cable, final attachment and electronic check of the data logger, deployment and latching of the thermistor assembly (confirmed with 5-kips wireline overpull), landing and latching of the CORK assembly (confirmed with 10-kips drill pipe overpull), recovery of the CORK setting go-devil, deployment of the CORK ROV platform, VIT deployment, release of the CORK running tool, and VIT recovery. The only negative factor to the final CORK operation was that the lower portion of the CORK setting go-devil backed off during recovery and was lost on the seafloor. While storing the CORK tools and securing the rig floor for transit, the four lower guide horn (LGH) cables (three, each broken) were removed from the moonpool area, so that new cables could be fabricated before arrival in Victoria. The upper guide horn (UGH) was installed and the ship got underway for the first (Site 1029) of several alternate sites to be cored during the remaining operational time.
After completion of Site 1026, the ship sailed for 2.5 hr to Site 1029. The BHA was made up and run to bottom for a routine APC/XCB cored hole. The driller tagged bottom while advancing to the APC shoot depth. Accordingly, the bit was picked up 5 m and Hole 1029A was spudded. The seafloor depth was established at 2664.0 mbrf, or 7.4 m above the PDR depth. Coring proceeded for 12 APC cores during which five Adara temperature measurements were taken. APC coring was terminated as a result of incomplete strokes and an overpull of 100,000 lbs was reached. The XCB system was used to deepen the hole with excellent core recovery until basement was tagged at a depth of 2883.5 mbrf (219.5 mbsf). The hole was terminated at 2887.0 mbrf (223.0 mbsf) after coring basement for 1 hr. The drill string was recovered to approximately 100 m above seafloor and the beacon was released/recovered. With knobby drilling joints installed through the guide horn area, the ship began moving to Site 1030.
Hole 1030A was located on a ridge crest with approximately 30 m of sediment cover overlying basement. A positioning beacon was deployed and the drill string was spaced out for initiating drilling operations. A wash test was conducted to avoid hitting basement with the APC and to allow optimal placement of the DVTP for a temperature measurement as close as possible to basement. With a core barrel in place, the drill string was lowered while circulating at a slow rate until the mudline was tagged at a depth of 2589.0 mbrf. Circulation and rotation were kept at a minimal rate until reaching a perceived basement depth of 2628.0 mbrf (39.0 mbsf). The drill string was pulled clear of the mudline and the vessel was offset 10 m for spudding Hole 1030A. The drill string was spaced out for shooting the first APC core from a depth of 2586.0 mbrf, or 3.0 m above tag depth. The spaceout was cut closer than normal, because the scientists desired a minimum 6.5 to 7.0 m of core recovery on the mudline core. This was to be used for chemistry analysis, which would determine the location of the next site. Two potential sites were under consideration after coring Site 1030, and the deciding factor was the chemical analysis from Hole 1030A. When the first core was recovered, it was full, indicating a missed mudline depth. The core was curated as a "wash" Core 168-1030A-1W and coring was halted after this single core. The pipe was then pulled clear of the mudline, ending Hole 1030A.
The bit was placed at 2579.0 mbrf, or 7.0 m higher for spudding Hole 1030B. No offset in DP mode was made between holes. This time, 4.01 m was recovered in the first core, establishing the mudline at 2584.5 mbrf. Continuous APC coring continued through Core 168-1030B-5H to a depth of 2625.5 mbrf (41.0 mbsf). The DVTP was then deployed for a single temperature measurement close to the top of anticipated basement. A single XCB Core 168-1030B-6X was then cored through the remaining sediment and into basement. The drillers basement tag depth was 2631.4 m (46.9 mbsf). The drill string was pulled clear of the sea bed and secured at a depth of 2469.4 mbrf (approximately 100 m above the seafloor) with knobby drilling joints through the guide horn. The beacon eventually released after multiple transmissions and was recovered aboard ship. This was the first and only beacon problem experienced during the leg. Once the beacon was secured, the ship began to move in DP mode to Hole 1031A approximately 1000 m away.
The DP move to Hole 1031A was uneventful. Once on location, a positioning beacon was deployed and the pipe was positioned to wash to basement in a manner similar to Site 1030. Jetting was initiated at 1135 hr and the pipe was advanced to a depth of 2642.5 mbrf (65.5 mbsf), where apparent basement was tagged. The pipe was pulled clear of the sea bed, and the ship was offset 10 m west for spudding Hole 1031A. Apparently, the PDR reading was off because of a side echo from a nearby seafloor feature. After two water cores, it was decided to advance the drill string and feel for bottom, and then pick back up for spudding. Two additional single joints were added to the string and seafloor was eventually tagged at ~2602 mbrf. The drill pipe was placed at 2598.5 mbrf, and a mudline core totaling 8.86 m was recovered. APC coring continued through Core 168-1031A-5H with Adara temperature measurements taken on Cores 168-1031A-3H and 4H. The DVTP was deployed after Core 168-1031A-4H. Core 168-1031A 5H impacted basement and recovered approximately 4.0 m of undisturbed core with the remainder of the material flow-in. A single XCB Core 168-1031A-6X was advanced 1.5 m into basement in just under 1 hr, recovering 0.38 m of basement rock. The drill string was pulled out of the hole and the APC/XCB BHA was nondestructively tested as it reached the drill floor. Once secured, the beacon was recovered and the vessel departed for Site 1032 (Proposed Site LH-2).
The transit to the last site took only 2.5 hr from getting under way to the beacon drop. Hole 1032A was spudded with a mudline tag depth of 2656.0 mbrf. Drilling, with a center bit in place, proceeded to a depth of 2840.5 mbrf (184.5 mbsf). The center bit was recovered and a single RCB core was cut before the first of six DVTP deployments. Continuous RCB coring, interspersed with DVTP measurements, continued to a depth of 2947.2 mbrf (291.2 mbsf), where basement was tagged 1 m in on Core 168-1032A-12R. Basement coring continued to a TD of 2994.4 mbrf (338.4 mbsf, "Site Summary table"). As the hole progressed farther into basement there was increasingly more fill between connections. Three 40-bbl sepiolite mud sweeps and one 40-bbl sweep with high-viscosity gel mud were made during basement coring to aid in cleaning the hole of cuttings. After coring was terminated, a short trip above the basement contact was made and 14 m of fill was identified on bottom. Another 40-bbl sweep was made while circulating to bottom through the soft fill. A wiper trip to 100 mbsf and back to TD was made prior to logging. Two additional 40-bbl sweeps of gel mud were made during the wiper trip and before displacing the hole with 123 bbl of sepiolite mud for logging. The bit was released and the MBR sleeve was reverse shifted. During the second run, the coring line was coated because it was the last wireline run of the leg. The pipe was pulled to a logging depth of 2746.3 mbrf (90.3 mbsf) with 20 and 30-ft knobbies installed. The logging program consisted of three suites. The first run was made with the triple-combination consisting of porosity, density, photoelectric effect, resistivity, and gamma ray (DITE/HLDT/APS/HNGS tools). In addition, the Lamont-Doherty temperature logging tool (TLT) was run. The first suite of logging tools reached a depth of 2938.0 mbrf (282.0 mbsf). This was a disappointing 9.2 m above the basement contact. Efforts were made to go deeper, but to no avail. The caliper logs recovered during this run indicated a fairly good hole through the sediment column (11"-12" hole) but a very ratty hole in the deeper part of the section above basement. The second suite of logs consisted of the FMS and sonic tools (FMS/sonic/NGT). This suite reached a depth of 2933.0 mbrf, or 5 m shallower than the first run. Two passes with the FMS/Sonic tools were made. The third logging suite consisted of the geochemical tool (GST/ACT/CNTG/NGTC). The third deployment reached a depth of 2932.0 mbrf, another meter higher in the hole. Logging was completed by 1900 hr, 14 August 1996. The drill pipe was pulled and the rig floor was secured at 2400 hr, 14 August. The ship set sail for Victoria, BC, with an estimated arrival at 2100 hr, 15 August.
To 168 Table of Contents